How would I solve this equation? I have to solve for y:
20 + 3/4(12+y) = 3/2y - 10
Please help me out! :) Also, if you can show the steps? Thanks a bunch!!

Answer by stanbon(75887)

20 + 3/4(12+y) = 3/2y - 10
---
Add 10 to both sides:
----
30 + (3/4)(12+y) = (3/2)y
----
30 + 9 + (3/4) = (6/4)y
====
39 = (3/4)y
---
13 = (1/4)y
y = 52
================
Cheers,
Stan H.
============